OpenStudy (anonymous):

solve the system of equations y = X^2-6x-2, 4y-x=1

OpenStudy (kropot72):

One way to solve is to use substitution. The equation 4y - x =1 can be rearranged to give x in terms of y as follows: Subtracting 4y from each side gives -x = 1 - 4y Multiplying both sides by -1 gives x = 4y - 1.................(1) Now the value for x in equation (1) can be substituted in the quadratic. Can you do that?

OpenStudy (anonymous):

yeah but when i substitute this in i get a answer that i cannot factor and that is where i get stuck

OpenStudy (anonymous):

quadratic formula always works

OpenStudy (kropot72):

Substituting for y did you get \[y=16y ^{2}-8y+1-24y+6-2\]

OpenStudy (anonymous):

yeah i did get that

OpenStudy (kropot72):

So now we can form a quadratic and solve it. As @satellite73 pointed out all quadratics can be solved (by the formula if necessary). \[16y ^{2}-33y+5=0\] Can you solve this using the formula?

OpenStudy (anonymous):

yes thank you :)

OpenStudy (kropot72):

You're welcome :)